OpenStudy (lalaland_lauren):

Joyce is trying to solve the equation y = x2 − 8x + 7 using the quadratic formula. She has made an error in one of the steps below. Find the step where Joyce went wrong. Step 1: x equals negative 8 plus or minus the square root of the quantity eight squared minus four times one times seven, end quantity, all over two times one. Step 2: x equals negative 8 plus or minus the square root of sixty-four minus twenty-eight all over two times one. Step 3: x equals negative 8 plus or minus the square root of thirty-six all over two times one. Step 4: x equals negative 8 plus or minus six all over

OpenStudy (anikhalder):

So, let's go through this. Do you know the quadratic formula and how you can use it to solve this equation?

OpenStudy (lalaland_lauren):

No :(

OpenStudy (anikhalder):

No problem, it goes like this: for an equation: \[ax ^{2}+bx+c = 0\] You can solve for x using the following formula (quadratic formula): \[x = \frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a}\] Hence, in the given equation y = x2 − 8x + 7 and when you set it to zero, i.e. \[x ^{2}− 8x + 7 = 0\] Then our a, b and c are as follows: a = 1 b = -8 c = 7 Now, just plug these into the quadratic formula and see which one out of the options you have is not satisfied. :)))

OpenStudy (lalaland_lauren):

Whoa that's a lot of letters Dx I'm sorry, I'm not exactly smart when it comes to these things

OpenStudy (anikhalder):

https://www.khanacademy.org/math/algebra/quadratics/quadratic-formula/v/using-the-quadratic-formula Try this video :))) Much better than my explanation :P

OpenStudy (lalaland_lauren):

Wait wait wait never mind! I think I got it now :)

OpenStudy (anikhalder):

Uber cool ! :)))

OpenStudy (jacob902):

Joyce is trying to solve the equation y = x2 − 8x + 7 using the quadratic formula. She has made an error in one of the steps below. Find the step where Joyce went wrong. Step 1: x equals negative 8 plus or minus the square root of the quantity eight squared minus four times one times seven, end quantity, all over two times one. Step 2: x equals negative 8 plus or minus the square root of sixty-four minus twenty-eight all over two times one. Step 3: x equals negative 8 plus or minus the square root of thirty-six all over two times one. Step 4: x equals negative 8 plus or minus six all over two. Step 1 Step 2 Step 3 Step 4

OpenStudy (jacob902):

Joyce is trying to solve the equation y = x2 − 8x + 7 using the quadratic formula. She has made an error in one of the steps below. Find the step where Joyce went wrong. FIRST OF ALL, Y MUST EQUAL 0 Step 1: x equals negative 8 plus or minus the square root of the quantity eight squared minus four times one times seven, end quantity, all over two times one. THIS IS WRONG. IT SHOULD BE POSITIVE 8 plus or minus the square root of the quantity eight squared minus four times one times seven, end quantity, all over two times one. Step 2: x equals POSITIVE 8 plus or minus the square root of sixty-four minus twenty-eight all over two times one. Step 3: x equals POSITIVE 8 plus or minus the square root of thirty-six all over two times one. Step 4: x equals POSITIVE 8 plus or minus six all over two. Step 5 X=(8±6)/2 Step 6 X=14/2 =7 or Step 7 X=2/2 =1 Answers {1,7}